Write down the line structures of the compound CH3CHC(OH)CH3
1
Expert's answer
2020-07-16T07:01:23-0400

The line structures of the compound CH3CHC(OH)CH3 are listed below:

CH3-CH2-CH2-HC=O

CH3-CO-CH2-CH3

СH2=CH-CH2-CH2OH

CH3-CH=CH-CH2OH

СН2=СН-СН (ОН)-СН3

СН2=СН-СН2-О-СН3